# Constrained Stochastic Recursive Momentum Successive Convex Approximation

### v2

@@ Appendix B Proof of Lemma 2
$\langle{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t},{\mathbf{z}}_{t+1}+{\mathbf{v}}_{{\hat{{\mathbf{x}}}}_{t}}\rangle+\langle{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t},\nabla{\tilde{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})-\nabla{\tilde{f}}({\mathbf{x}}_{t},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})\rangle\leq 0.$ (76)
@@
${\mathbb{E}}\left[F({\mathbf{x}}_{t+1})-F({\mathbf{x}}_{t})\right]\leq\frac{1}{2}{\mathbb{E}}\left[\eta_{t}\left\|\nabla U({\mathbf{x}}_{t})-{\mathbf{z}}_{t+1}\right\|^{2}\right]$
$~{}~{}~{}+\frac{L}{2}{\mathbb{E}}\left[\eta_{t}^{2}\left\|{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t}\right\|^{2}\right]+\bigg{(}\frac{1}{2}-\mu\bigg{)}{\mathbb{E}}\left[\eta_{t}\left\|{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t}\right\|^{2}\right]$
@@ Appendix C Proof of Lemma 4
Since ${\mathbf{x}}_{t}\in{\mathcal{X}}$, strong MFCQ holds at ${\mathbf{x}}_{t}$, implying the existence of ${\mathbf{d}}_{t}$ as specified in Assumption (A5). Consider the point ${\tilde{\mathbf{x}}}({\mathbf{x}}_{t})={\mathbf{x}}_{t}+\gamma{\mathbf{d}}_{t}$ for $\gamma>0$ and let ${\bar{\mathcal{I}}}$ and ${\bar{\mathcal{J}}}$ be defined as the sets of indices corresponding to the near-active constraints at ${\mathbf{x}}_{t}$ for the convex and general constraints, respectively.
@@ Appendix D Proof of Lemma 5
${\tilde{f}}({\mathbf{x}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})+u({\mathbf{x}})$ $\geq{\tilde{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})+u({{\hat{{\mathbf{x}}}}_{t}})+\langle\nabla{\tilde{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})+{\mathbf{v}}_{{{\hat{{\mathbf{x}}}}_{t}}},{\mathbf{x}}-{{\hat{{\mathbf{x}}}}_{t}}\rangle$ (91)
@@
${\tilde{f}}({\tilde{\mathbf{x}}}({\mathbf{x}}_{t}),{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})+u({\tilde{\mathbf{x}}}({\mathbf{x}}_{t}))+\sum\nolimits_{j=1}^{J}[{\hat{{\boldsymbol{\lambda}}}}_{t}]_{j}{\tilde{g}}_{j}({\tilde{\mathbf{x}}}({\mathbf{x}}_{t}),{\mathbf{x}}_{t})+\sum\nolimits_{i=1}^{I}\nu_{i}({{\hat{{\mathbf{x}}}}_{t}})h_{i}({\tilde{\mathbf{x}}}({\mathbf{x}}_{t}))$
$\geq{\tilde{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})+u({{\hat{{\mathbf{x}}}}_{t}})+\sum\nolimits_{j=1}^{J}[{\hat{{\boldsymbol{\lambda}}}}_{t}]_{j}{\tilde{g}}_{j}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t})+\sum\nolimits_{i=1}^{I}\nu_{i}({{\hat{{\mathbf{x}}}}_{t}})h_{i}({{\hat{{\mathbf{x}}}}_{t}})$
$+\Big{\langle}\sum\nolimits_{j=1}^{J}[{\hat{{\boldsymbol{\lambda}}}}_{t}]_{j}\nabla{\tilde{g}}_{j}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t})+\sum\nolimits_{i=1}^{I}\nu_{i}({{\hat{{\mathbf{x}}}}_{t}})\nabla h_{i}({{\hat{{\mathbf{x}}}}_{t}})+\nabla{\tilde{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})+{\mathbf{v}}_{{{\hat{{\mathbf{x}}}}_{t}}},{\tilde{\mathbf{x}}}({\mathbf{x}}_{t})-{{\hat{{\mathbf{x}}}}_{t}}\Big{\rangle}$ (94)
@@
$\geq-\sum\nolimits_{j=1}^{J}[{\hat{{\boldsymbol{\lambda}}}}_{t}]_{j}{\tilde{g}}_{j}({\tilde{\mathbf{x}}}({\mathbf{x}}_{t}),{\mathbf{x}}_{t})-\sum\nolimits_{i=1}^{I}\nu_{i}({{\hat{{\mathbf{x}}}}_{t}})h_{i}({\tilde{\mathbf{x}}}({\mathbf{x}}_{t}))$ (95)
$\geq\frac{\rho^{2}}{2L}\bigg{(}\sum_{j=1}^{J}[{\hat{{\boldsymbol{\lambda}}}}_{t}]_{j}+\sum_{i=1}^{I}\nu_{i}({{\hat{{\mathbf{x}}}}_{t}})\bigg{)}.$ (96)

### v3

@@ Appendix B Proof of Lemma 2
$\langle{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t},{\mathbf{z}}_{t+1}+{\mathbf{v}}_{{\hat{{\mathbf{x}}}}_{t}}\rangle$ $~{}+\langle{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t},\nabla{\tilde{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})-\nabla{\tilde{f}}({\mathbf{x}}_{t},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})\rangle\leq 0.$ (82)
@@
${\mathbb{E}}\left[F({\mathbf{x}}_{t+1})-F({\mathbf{x}}_{t})\right]$ $\leq\frac{1}{2}{\mathbb{E}}\left[\eta_{t}\left\|\nabla U({\mathbf{x}}_{t})-{\mathbf{z}}_{t+1}\right\|^{2}\right]+\frac{L}{2}{\mathbb{E}}\left[\eta_{t}^{2}\left\|{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t}\right\|^{2}\right]+\bigg{(}\frac{1}{2}-\mu\bigg{)}{\mathbb{E}}\left[\eta_{t}\left\|{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t}\right\|^{2}\right]$
@@ Appendix C Proof of Lemma 4
Since ${\mathbf{x}}_{t}\in{\mathcal{X}}$, parameterized MFCQ holds at ${\mathbf{x}}_{t}$, implying the existence of ${\mathbf{d}}_{t}$ as specified in Assumption (A7). Consider the point ${\tilde{\mathbf{x}}}({\mathbf{x}}_{t})={\mathbf{x}}_{t}+\gamma{\mathbf{d}}_{t}$ for $\gamma>0$ and let ${\bar{\mathcal{I}}}$ and ${\bar{\mathcal{J}}}$ be defined as the sets of indices corresponding to the near-active constraints at ${\mathbf{x}}_{t}$ for the convex and general constraints, respectively.
@@ Appendix D Proof of Lemma 5
${\tilde{f}}({\mathbf{x}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})7+u({\mathbf{x}})$ $\geq{\tilde{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})+u({{\hat{{\mathbf{x}}}}_{t}})+\langle\nabla{\tilde{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})+{\mathbf{v}}_{{{\hat{{\mathbf{x}}}}_{t}}},{\mathbf{x}}-{{\hat{{\mathbf{x}}}}_{t}}\rangle$ (97)
@@
${\tilde{f}}({\tilde{\mathbf{x}}}({\mathbf{x}}_{t}),{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})+u({\tilde{\mathbf{x}}}({\mathbf{x}}_{t}))+\sum\nolimits_{j=1}^{J}[{\hat{{\boldsymbol{\lambda}}}}_{t}]_{j}{\tilde{g}}_{j}({\tilde{\mathbf{x}}}({\mathbf{x}}_{t}),{\mathbf{x}}_{t})+\sum\nolimits_{i=1}^{I}[\hat{\nu}_{t}]_{i}h_{i}({\tilde{\mathbf{x}}}({\mathbf{x}}_{t}))$
$\geq{\tilde{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})+u({{\hat{{\mathbf{x}}}}_{t}})+\sum\nolimits_{j=1}^{J}[{\hat{{\boldsymbol{\lambda}}}}_{t}]_{j}{\tilde{g}}_{j}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t})+\sum\nolimits_{i=1}^{I}[\hat{\nu}_{t}]_{i}h_{i}({{\hat{{\mathbf{x}}}}_{t}})$
$+\Big{\langle}\sum\nolimits_{j=1}^{J}[{\hat{{\boldsymbol{\lambda}}}}_{t}]_{j}\nabla{\tilde{g}}_{j}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t})+\sum\nolimits_{i=1}^{I}[\hat{\nu}_{t}]_{i}\nabla h_{i}({{\hat{{\mathbf{x}}}}_{t}})+\nabla{\tilde{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})+{\mathbf{v}}_{{{\hat{{\mathbf{x}}}}_{t}}},{\tilde{\mathbf{x}}}({\mathbf{x}}_{t})-{{\hat{{\mathbf{x}}}}_{t}}\Big{\rangle}$ (100)
@@
$\geq-\sum\nolimits_{j=1}^{J}[{\hat{{\boldsymbol{\lambda}}}}_{t}]_{j}{\tilde{g}}_{j}({\tilde{\mathbf{x}}}({\mathbf{x}}_{t}),{\mathbf{x}}_{t})-\sum\nolimits_{i=1}^{I}[\hat{\nu}_{t}]_{i}h_{i}({\tilde{\mathbf{x}}}({\mathbf{x}}_{t}))$ (101)
$\geq\frac{\rho^{2}}{2L}\bigg{(}\sum_{j=1}^{J}[{\hat{{\boldsymbol{\lambda}}}}_{t}]_{j}+\sum_{i=1}^{I}[\hat{\nu}_{t}]_{i}\bigg{)}.$ (102)
@@
Appendix E Proof of Corollary 2 and Corollary 3
Proof:
Under non-adaptive step sizes $\eta_{t}=\eta=\tfrac{{\bar{k}}}{T^{1/3}}$ and $\beta_{t}=\beta=c\eta^{2}$, the modified results in Lemmas 1 and 2 stated as follows,
$\frac{1}{16L^{2}\eta}{\mathbb{E}}\left[\left\|{{\boldsymbol{\varepsilon}}}_{t+1}\right\|^{2}\right]\leq\frac{(1-\beta)}{16L^{2}\eta}{\mathbb{E}}\left[\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|^{2}\right]+\frac{\eta}{8}{\mathbb{E}}\left[\left\|{\boldsymbol{\delta}}_{t}\right\|^{2}\right]+\frac{\beta^{2}}{8L^{2}\eta}{\mathbb{E}}\left[(G_{t+1})^{2}\right],$ (103)
and
${\mathbb{E}}\left[F({\mathbf{x}}_{t+1})-F({\mathbf{x}}_{t})\right]\leq\frac{\eta}{2}{\mathbb{E}}\left[\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|^{2}\right]-\frac{\eta}{4}{\mathbb{E}}\left[\left\|{\boldsymbol{\delta}}_{t}\right\|^{2}\right].$ (104)
By introducing a Lyapunov function $\Psi_{t}={\mathbb{E}}\left[F({\mathbf{x}}_{t})-F^{\star}\right]+\tfrac{1}{16L^{2}\eta}{\mathbb{E}}[\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|^{2}]$ and adding (103) with (104), we obtain
$\Psi_{t+1}-\Psi_{t}$ $\leq\frac{\alpha}{16L^{2}}{\mathbb{E}}\left[\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|^{2}\right]-\frac{\eta}{8}{\mathbb{E}}\left[\left\|{\boldsymbol{\delta}}_{t}\right\|^{2}\right]+\frac{c^{2}\eta^{3}}{8L^{2}}{\mathbb{E}}\left[G_{t+1}^{2}\right],$ (105)
with $\alpha=8L^{2}\eta+\frac{(1-\beta)}{\eta}-\frac{1}{\eta}$. Taking telescopic sum from $t=1\ldots T$,
$\Psi_{T+1}-\Psi_{1}$ $\leq\frac{\alpha}{16L^{2}}\sum_{t=1}^{T}{\mathbb{E}}\left[\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|^{2}\right]-\frac{\eta}{8}\sum_{t=1}^{T}{\mathbb{E}}\left[\left\|{\boldsymbol{\delta}}_{t}\right\|^{2}\right]+\frac{c^{2}\eta^{3}}{8L^{2}}\sum_{t=1}^{T}G^{2}$ (106)
Also, we have that
$\alpha$ $=8L^{2}\eta-\frac{\beta}{\eta}=-2L^{2}\eta\left(\frac{c}{2L^{2}}-4\right)=-2L^{2}d^{\prime}\eta$ (107)
where $d^{\prime}:=\frac{c}{2L^{2}}-4>0$. Here, since ${\mathbb{E}}[\left\|{{\boldsymbol{\varepsilon}}}_{1}\right\|^{2}]\leq\sigma^{2}$ from Assumption A5, we can bound
$\Psi_{1}\leq F({\mathbf{x}}_{1})-F^{\star}+\frac{\sigma^{2}}{16L^{2}\eta_{0}}=B_{1}+\frac{\sigma^{2}}{16L^{2}{\bar{k}}}.$ (108)
Rearranging (105) we can write,
$\tfrac{1}{T}\sum_{t=1}^{T}{\mathbb{E}}\left[\left\|{\boldsymbol{\delta}}_{t}\right\|^{2}\right]\leq\frac{M^{\prime}_{T}}{T^{2/3}}={\mathcal{O}}\left(\frac{1}{T^{2/3}}\right)$ (109)
where $M^{\prime}_{T}:=8\frac{B_{1}}{{\bar{k}}}+\frac{\sigma^{2}}{2L^{2}\bar{k}^{2}}+\frac{c^{2}\bar{k}^{2}}{L^{2}}G^{2}$. Similarly,
$\tfrac{1}{T}\sum_{t=1}^{T}{\mathbb{E}}\left[\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|^{2}\right]\leq\frac{\bar{M^{\prime}}_{T}}{d^{\prime}T^{2/3}}={\mathcal{O}}\left(\frac{1}{T^{2/3}}\right)$ (110)
where $d^{\prime}=\tfrac{1}{2L^{2}}(c-8L^{2})$. Finally, using similar inequalities as (41) and (42) we obtain the results of Corollary 2.
Also, under non-adaptive step size, in light of Corollary 2, (60) can be bounded as,
$\tfrac{1}{T}\sum\nolimits_{t=1}^{T}{\mathbb{E}}\left[\left\|{\boldsymbol{\pi}}_{t}\right\|\right]^{2}$ $\leq\tfrac{2}{T}\sum_{t=1}^{T}\left(2L+\frac{4B_{U}L^{2}}{\rho^{2}}\right)^{2}{\mathbb{E}}\left[\left\|{\boldsymbol{\delta}}_{t}\right\|\right]^{2}+\tfrac{2}{T}\sum_{t=1}^{T}{\mathbb{E}}\left[\left\|{\boldsymbol{\varepsilon}}_{t}\right\|\right]^{2}$
$\leq{\mathcal{O}}(T^{-2/3})$ (111a)
Similarly, in light of Corollary 2, (63) can bounded as,
$\tfrac{1}{T}\sum\nolimits_{t=1}^{T}{\mathbb{E}}[{\hat{{\boldsymbol{\lambda}}}}_{t}^{\mathsf{T}}{\mathbf{g}}({{\hat{{\mathbf{x}}}}_{t}})]$ $\geq-\frac{2B_{U}L^{2}}{\rho^{2}}{\mathbb{E}}[\left\|{\boldsymbol{\delta}}_{t}\right\|^{2}]={\mathcal{O}}(T^{-2/3}).$ (111b)
Therefore,
$\min_{1\leq t\leq T}({\mathbb{E}}\left[\left\|{\boldsymbol{\pi}}_{t}\right\|\right]^{2}-{\mathbb{E}}[{\hat{{\boldsymbol{\lambda}}}}_{t}^{\mathsf{T}}{\mathbf{g}}({{\hat{{\mathbf{x}}}}_{t}})])$ $\leq\tfrac{1}{T}\sum\nolimits_{t=1}^{T}{\mathbb{E}}\left[\left\|{\boldsymbol{\pi}}_{t}\right\|\right]^{2}-\tfrac{1}{T}\sum\nolimits_{t=1}^{T}{\mathbb{E}}[{\hat{{\boldsymbol{\lambda}}}}_{t}^{\mathsf{T}}{\mathbf{g}}({{\hat{{\mathbf{x}}}}_{t}})]$
$\leq{\mathcal{O}}(T^{-2/3})$ (111c)
Hence, and we obtain the desired result in Corollary 3